Grade 12th passAlgebra

A rational number is such that when you multiply it by 5/2 and add 2/3 to the product, you get -7/12. What is the number?

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

Let the rational be x.
According to the question,
x×(5/2)+2/3=-7/12
⇒5x/2+2/3=-7/12
⇒5x/2=-7/12–2/3
⇒5x/2 = (-7- 8)/12
⇒5x/2 = -15/12
⇒5x/2 = -5/4
⇒x=(-5/4)×(2/5)
⇒x=-10/20
⇒x=-1⁄2
Therefore, the rational number is -1⁄2.

Thanks